OpenStudy (anonymous):

I need help on a math question. The question says, "Jeremy had 3/4 of a submarine sandwich and gave his friend 1/3 of it. What fraction of the sandwich did the friend receive?" Do I divide and use the reciprocal to multiply it then simplify?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

one third of three fourths is \[\huge \frac{1}{3}\times \frac{3}{4}\]

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So, is that my answer?

OpenStudy (anonymous):

yup